SAT For a specific year, the national average score on the SAT Math test was 515. 496 A superintendent wishes to see if their493 below those cells. 506 551 500 Lower Upper 504iPad Kindle 9 10 Consumer Reports would like to conduct a hypothesis test to determine if the average battery life for an iPa

Please help me with what formulas to put in to get the right answers. I have tried but I am stuck not getting it... For the first one I am only able to find a two paired t-test but there is only one row so I don't know what to do...
Answer #1

I) a) a t-test allows us to compare the average values of the two data sets and determine if they came from the same population.

using t test here we wat to test mu=515 against mu=! 515

our test statistics is

X – μ t= Z — S ô / vn

b)confidence interval is Lower limit = M - Z.95σM

Upper limit = M + Z.95σM

where M is the mean

II) a)The two-sample t-test for unpaired data is defined as: Ho: Mi = 42 Mi + u2 Test Statistic: Y-Y, s/Ni+s/N, Hai T= where Ny and

b)